.


 

A shipment of 9 printers contains 3 that are defective. Find the probability that a sample of size 3​, drawn from the 9​, will not contain a defective printer.


1
Expert's answer
2021-04-20T16:26:37-0400

3 printers out of 10 can be chosen "n = C_9^3" ways - total number of outcomes.

The number of favorable outcomes is "m = C_{9 - 3}^3 = C_6^3".

Then the wanted probability is

"p = \frac{m}{n} = \frac{{C_6^3}}{{C_9^3}} = \frac{{6!}}{{3!3!}} \cdot \frac{{3!6!}}{{9!}} = \frac{{4 \cdot 5 \cdot 6}}{{7 \cdot 8 \cdot 9}} = \frac{5}{{7 \cdot 3}} = \frac{5}{{21}}"

Answer: "p = \frac{5}{{21}}"